# SODAs: Sparse Optimization for the Discovery of Differential and Algebraic Equations

### v1

@@ 4.2 Example 2: Non-linear and Chaotic Pendulums
    We used a Pareto front of the log of the condition number of the library to determine the stopping criterion. In Case 3, we stopped after two iterations as no further improvement in the condition number was observed from the third refinement onward (Figure 4d), implying the existence of two algebraic constraints. We also conducted a comparative study of the data requirement for Cases 2 and 3 as a function of maximum degree of the monomials in the candidate library (Figure 4c). The results indicate that the data requirement, in terms of the number of frames per video segment, stayed within manageable limits. Interestingly, the double pendulum required fewer frames per video segment for successful recovery of the algebraic equations compared to Case 2. This is because the chaotic nature of the pendulum, combined with a 3-minute-long video, provided a good sampling of the state space. We also shuffled the data points along the time axis only for Case 3 to ensure a good representation of the dynamics, given the longer length of the video segment. These factors contributed to the reduced data requirement in terms of the number of frames needed for successful recovery in Case 3. For Case 2, as the degree of the library reaches 5, the data requirement increases sharply. This can attributed to the fact that higher-order terms in the library tend to be highly correlated when there is limited sampling of the state space. An alternative approach to reduce the number of frames required would be to conduct experiments with more informative initial conditions, which is closely related to experimental design [65, 66].
    4.3 Example 3: Power-grid Dynamics
    The operation of modern power grids relies on detailed modeling of dynamical behavior and network topology, which are critical for grid stability, efficiency, and control against disturbances [67]. Such model information directly supports daily operations, scheduling, and advanced techniques like state estimation, fault monitoring, and optimal power flow computation. However, the rise of renewable generation has led to decentralized power production, making detailed grid models less accessible in distribution systems due to frequent structural modifications. These include the addition and removal of network components and generating units, high uncertainty in distributed energy resources, line maintenance, faults, and other unplanned events. Yet, traditional power-grid monitoring and fault detection methods are often grounded on parameter estimation techniques, requiring extensive prior knowledge of grid configurations and physical properties [68, 69, 70, 71]. The increasing deployment of phasor measurement units (PMUs) [72], capable of providing high-resolution data, offers a significant opportunity for automated model discovery [73, 74]. Here, we show that the sparse discovery of differential-algebraic models infers network topology robustly, which is encoded by the algebraic relationships among power levels at each bus. Using numerical results, we demonstrate that the proposed method accurately infers network models across a wide range of scenarios, including both small- and large-scale power grids, varying levels of state perturbation, and diverse PMU signal qualities.
    System description.
    The dynamics of power grids can be modeled as a network of coupled oscillators (nodes) interconnected through power transmission lines (links) [75, 76, 77]. The nodes correspond to conventional generators and consumers (geographically represented by substations and their attached load). In this representation, the dynamics of generators are governed by the swing equation [78]:
    SODAs application.
    We considered three IEEE benchmark cases with varying complexity (Figure 5a): IEEE-4, IEEE-9, and IEEE-39. Time series were generated under two distinct conditions: i) strong perturbations (Figure 5b - top), which excite nonlinear transient behaviors, and ii) weak perturbations (Figure 5b - bottom), suitable for small-signal stability analysis. We assume that the power grid is equipped with phasor measurement units (PMUs) placed on all $N$ buses, providing us with high-resolution, high-frequency data for phase angle $\phi_{i}(t)$, frequency $\dot{\phi}_{i}(t)$, line voltage $V_{i}(t)$, and electrical power $P_{\rm e,i}(t)$ and $Q_{\rm e,i}(t)$. More information on network details, simulations and practicality of assumptions can be found in the supplementary Section 1.3.
    SODAs was applied to the simulated time series to identify the network structure and parameters of the governing DAE system as described in equations (36)–(38). By using the well-established physical principles of power-grid dynamics, we achieved a significant advantage in constraining SODAs within a more informed framework. More specifically, the conservation of active power at each node leads to inherent algebraic constraints that we exploited to improve the identification process. In constructing the candidate library, we included the following functional forms: $P_{{\rm e},i}$ (active power states), $\phi_{i}$ (phase angles), $\omega_{i}$ (frequency), and sine terms of phase differences $\sin(\phi_{i}-\phi_{j})$ for all pairs $(i,j)$. In a network with $N$ nodes, this initial library included $3N+\frac{N(N+1)}{2}$ terms, which amounts to 39, 114, and 1,372 terms for the IEEE-4, IEEE-9, and IEEE-39 benchmarks, respectively. Additionally, by incorporating prior knowledge, we further reduced the number of candidate terms. For example, the conservation law involving $P_{{\rm e},i}$ at a given node $i$ restricts the algebraic relations to include only sine terms of the form $\sin(\phi_{i}-\phi_{j})$, for all nodes $j$. Using the Daefiender.FeatureCouplingTransformer class, we efficiently enforced these additional library constraints, minimizing computational complexity and enhancing the accuracy of identified models by eliminating irrelevant terms.
    In the algebraic finder step, power laws constraining the active powers $P_{{\rm e},i}$ for all $i$ are discovered (38). These power laws inherently reveal the power grid’s network structure (encoded by the admittance matrix $Y$). In the dynamic finder step of SODAs, the differential equations governing the generators are discovered using a SINDy framework. We approximated the left-hand side of the dynamic equations (36)–(37) by utilizing derivatives of spline-fitted phase angles and applied sparse regression techniques, where LASSO was followed by sequential thresholding. Parameter estimation was subsequently performed through linear fitting with the discovered model structure. Note that for the identification of the dynamic equation, we modified the candidate library by removing the $\sin(\phi_{i}-\phi_{j})$ terms, which had already been accounted for by the algebraic constraints. Instead, we incorporated state measurements of $P_{{\rm e},i}$ and hence reducing the number of candidate functions in the dynamic finder step from $3N+\frac{N(N+1)}{2}$ to $3N$. This refinement has also helped to avoid using the coupling terms $\sin(\phi_{i}-\phi_{j})$, which have high sensitivity to noise in the data, for the dynamic finder step. We observed complete recovery of the true ODE system whenever the algebraic constraints were discovered with 100% accuracy. In contrast, attempts to identify coupled ODE equations without segregating the algebraic terms led to poor model recovery , even at SNR as high as 40dB, emphasizing the importance of this methodological separation of algebraic and differential equations.
    We present SODAs’ performance on identifying the ground-truth model for all three test cases, with a primary focus on the IEEE-39 benchmark test case (Figure 5c and 5d). We conducted similar analyses on the IEEE-4 and IEEE-9 benchmark networks, where SODAs demonstrated performance comparable or superior to that of the more complex IEEE-39 case. To evaluate the robustness of SODAs, we introduced varying amounts of noise into the simulated time-series data and examined performance across four different signal-to-noise ratios (SNR): no noise, 40 dB, 30 dB, and 20 dB. For each SNR level, we generated five distinct noise realizations and analyzed the performance as an average across these realizations. We found no significant variations in performance among the different noise realizations within the same SNR level, which allowed us to exclude error bars related to noise realizations in our figures.
    For SNRs of no noise, 40 dB, and 30 dB, SODAs achieved 100% recovery of all algebraic relationships, as well as the dynamic ODEs, using the refined library for both small and large perturbations. Even under high noise conditions (SNR of 20 dB), SODAs maintained a recovery rate exceeding 80% for algebraic relations with 100 perturbations. The algebraic relationships related to power conservation at generator nodes were identified with relatively small number of perturbations, underscoring the significant role of these nodes in influencing the overall dynamics of the power grid. Techniques such as signal averaging and advanced filtering can improve the SNR, even when the inherent measurement SNR is lower. Therefore, an SNR of 30 dB is assumed to be realistic for our further analysis, as other studies have reported PMU devices with signal quality at or above this level [79, 80, 81, 82]. We evaluated the performance of SODAs at a 30 dB SNR across various benchmark cases and levels of perturbations (Figure 5d). We did a comparative analysis of the three benchmark cases, in terms of the number of perturbations required to recover all true DAE system components, and found that the number of perturbations required for full recovery increased with the complexity of the power grid, yet it remained within the realistic number of permutations observed in practice (Figure 5d - left). Also, strong perturbations that induce significant transients in system dynamics enabled quicker recovery of the DAE system as expected (Figure 5d - right).
    Figure 5: Discovery of power-grid networks. (a) Power-grid networks of the IEEE-4 (top left), IEEE-9 (bottom left), and IEEE-39 (right) benchmark systems. Generator buses are represented by blue nodes on the outer circle, while aggregated load buses are represented by gray nodes. (b) Time series of phase dynamics $\phi_{i}(t)$ for all nodes in the IEEE-39 power grid under different levels of perturbations: large perturbations (top), large perturbations zoomed-in (middle), and small perturbations zoomed-in (bottom). (c) Performance comparison of SODAs applied to IEEE-39 benchmark system across different SNR levels and : small perturbations (left) and large perturbations (right). (d) Performance comparison of SODAs at 30 dB SNR across: different benchmark cases (left) and different perturbation strengths (right).
@@ 5 Discussion
    There are several limitations and challenges worth pursuing in future work. One significant limitation of SODAs is the necessity to observe or measure all the state variables. In many physical systems, measuring all state variables can be challenging due to cost or physical infeasibility. Previous methods such as implicit-SINDy [26] and PI-SINDy [27] required no measurements of the algebraic states, but also assumed that all states were differential variables. Extending SODAs to discover DAEs even in the presence of unobserved states, potentially through recent strategies [19, 20, 21], would enable the identification of algebraic and differential variables and their equations when only a subset of states are measured. However, recovery in this setting will be even more challenging due to the non-convexity of sparse-optimization with hidden states as well as the potential lack of analytical identifiability in nonlinear DAE models [83].
@@
    One aspect of DAE discovery we did not yet explore is how SODAs performs when states are approximately but not fully in quasi-steady state. In realistic experimental sampling of systems with fast-slow variables, some transient dynamics may be sampled before the fast variables reach quasi-steady state. In this work, we used the reduced ODE formulation (24) of our CRNs to simulate the time series used for equation re-discovery, ensuring that the slow-variables were in perfect quasi-steady state. There was no error in the algebraic relationships since the algebraic variables were sampled during sufficiently short time interval of the transient regime. Future work is needed to investigate how SODAs performs as a function of sampling transient vs quasi-steady-state regimes for fast variables. An interesting avenue would be in mimicking the hybrid-SINDY approach [84, 85] in which a different model is selected when variables are in quasi-steady state.
@@
    There is a more general data limitation challenge inherent to sampling dynamical systems: sampling is often determined by the trajectories of the dynamics themselves which induces correlation between library terms. In cases where the available data does not adequately differentiate between lower-order and higher-order terms, strong correlations separate from the true algebraic relationships can occur. In Example 2, we showed how the data requirement increases as the degree of the monomial terms in the library increases to 5. This is attributed to the high correlation between higher-degree monomials in the library. This high correlation between library terms makes it challenging to separately identify their contributions to the dynamics of the system. This issue is not exclusive to SODAs; it is an inherent challenge in all SINDy-type methods that rely on sparse regression with function libraries. This challenge can be overcome by implementing improved data collection strategies or experimental design techniques [65, 66] that help with distinguishing different candidate library terms. We aim to develop methods to address these challenges in our future work. Finally, this study has focused on DAEs involving only ordinary derivatives, an exciting avenue for future research will be to extend SODAs to DAEs involving partial derivatives.
    Data Accessibility.
    All code and data used for numerical experiments in this paper are available at https://github.com/mjayadharan/DAE-FINDER_dev/tree/main/Examples. The DaeFinder package is also available through Python Package Index (PyPI) https://pypi.org/project/DaeFinder.
    Authors’ contributions.
@@
    M.J. developed the SODAs algorithm, implemented the DaeFinder package, and conducted all experiments (except for Example 1). C.C. contributed to the DaeFinder package and designed and conducted the numerical experiments in Example 1. A.N.M. designed the test cases and assisted in the numerical experiments in Example 3. N.M.M. designed and supervised the research. All authors edited the manuscript and approved the final draft for publication.
    Competing Interests.
    We have no competing interests.
    Funding.
@@
    This material is based on work supported by several funding agencies. N. M. M. and M. J. were supported by the U.S. Department of Energy, Office of Science, Office of Advanced Scientific Computing Research, under Award Number DE-SC0024253 and Army Contracting Command under Award Number W52P1J-21-9-3023. C. C. was supported by the National Science Foundation under Grant No. DGS-2021900. Any opinions, findings, and conclusions or recommendations expressed in this material are those of the author(s) and do not necessarily reflect the views of the Department of Energy, Army Contracting Command, or National Science Foundation.
    Acknowledgments.
@@
    We thank Finn Hagerty for his assistance in running simulations for the CRN example and related discussions.
    Supplementary Material
    1 Time series generation and Data processing
    1.1 Example 1: Chemical Reaction Networks
    Time series for three different CRNs were generated by solving equations of the simplified form (24). Note that this means that the algebraic relationships induced by quasi-steady-state approximations are strictly enforced. CRN1 is the simple Michealis-Menten reaction, CRN2 is the competition of two enzymes for a shared substrate, and CRN3 is the simplified propanediol utilization pathway. These reactions are:
    1. 1.
      CRN1: 1 reaction, 4 statesincluding complex and enzyme
    2. 2.
      CRN2: 2 reactions, 7 states including complexes and enzymes
    3. 3.
      CRN3: 4 reactions (one reversible), 14 states including complexes and enzymes
    CRN1 and CRN2 are nested within CRN3, with CRN3 chaining CRN1 and a modified CRN2 with an additional reaction consuming one of CRN2’s products. CRN3 also complicates the simple, irreversible reaction mechanism of CRN1 and CRN2, and one reaction is made reversible to mimic the current understandings of the true priopandiol utilization pathway. In function, reversibility can be thought of as having the effect of an additional reaction.
    Simulations for each CRN consisted of five initial conditions, each simulated to a fixed end time ($t=4$ for CRN1 & CRN2, $t=10$ for CRN3; Figure 3a). Within each CRN, initial enzyme concentrations were fixed between simulations, while initial substrate concentrations were selected to be equally spaced within a given range for each species (fixed for CRN1 & CRN2, set by reasonable experimental ranges for CRN3). Such a selection was intended to effectively sample the breadth of system behaviors that result from all possible initial conditions. Model parameters for CRN1 and CRN2 were chosen to meet the assumptions required for the quasi-steady-state approximation made in deriving Michaelis-Menten kinetics to be reasonable (i.e., that binding is much faster than unbinding and that initial substrate concentration far exceeds enzyme concentration) [37, 38]. Parameters for CRN3 were taken from the literature [57].
    Figure 6: CRN simulated time series: Five initial conditions simulated for each CRN at 1% noise with 100 data points per initial condition. Colored by species. (a) CRN1, (b) CRN2, (c) CRN3.
    To mimic possible measurement error present in real-world data and to test noise robustness of the SODAs algorithm, Gaussian noise of up to 15% was added to the simulated time series according to (20) prior to the application of SODAs. In addition, we applied SODAs across a varied sampling resolution, simulating uniform grids in $t$ with 5-1200 data points per initial condition for all CRNs. This is equivalent to $\Delta t\in\{0.00\bar{3},0.8\}$ for CRN1 and CRN2, and to $\Delta t\in\{0.008\bar{3},2\}$ for CRN3. Examples of the simulated time series from each CRN used to assess the SODAs algorithm are shown in Figure 6.
    1.2 Example 2: Non-linear and Chaotic Pendulum
    1. 1.
      Single pendulum experiments with negligible damping: Three separate experiments were conducted in which a single pendulum was swung from different initial positions and with varying initial velocities (Figure 4b - left). The initial conditions were chosen to ensure that small-angle or linear approximations were no longer valid. The experiments were recorded at 120 frames per second (fps), for a duration of 20 seconds. The recordings were made when the damping was minimal.
    2. 2.
      Heavily damped pendulum: Four different video clips of animation of a heavily damped pendulum were recorded at 60 fps, each lasting 20 seconds (Figure 7). Each clip captured various behaviors based on different initial conditions.
    3. 3.
      Chaotic double pendulum: A single video recording of the animation of a double pendulum (Figure 4b - right) exhibiting chaotic behavior at 60 fps was used. The recording lasted for 3 minutes. Both pendulums have equal length and the same masses attached to the end. Due to the chaotic and non-repetitive behavior of the system, this single initial condition provided a good sampling of various regions in the phase space.
    In all three cases, the OpenCV [86] package in Python was used to extract the pixel data from the video recordings. A combination of color- and motion-based sensing was used to track the masses of the pendulums and obtain their Cartesian coordinates. The raw pixel data was centered and scaled such that the origin was positioned at the pivot of the single pendulum and at the pivot of the primary pendulum in the double pendulum case. The scaling ensured that the length of the pendulum is 1. These preprocessing steps were performed to normalize the features and minimize unnecessary cross-terms in the discovered equations, even though we verified that SODAs could discover the correct equations without centering and scaling the pixel data. The raw footage and processed data sets used in this example are available at https://github.com/mjayadharan/DAE-FINDER_dev/tree/main/Examples/Example_2_pendulum/data.
    Figure 7: Damped pendulum: scaled pixel data.
    1.3 Example 3: Power-grid Dynamics
    We consider the following cases available through MATPOWER toolbox [87)]:
    1. 1.
      IEEE-4 benchmark: 2 generators and 4 buses.
    2. 2.
      IEEE-9 benchmark: 3 generators and 9 buses.
    3. 3.
      IEEE-39 benchmark: 10 generators and 39 buses.
    The IEEE-39 is a simplified representation of the New England section of the Eastern Interconnection in the North American power grid, where a single generator represents the link to the remainder of the grid. Each generator is associated to a terminal, represented as a white node, which acts as an interconnection point to the rest of the grid. The simulations include the nominal voltage magnitudes $V_{i}$, load consumption $P_{\ell,i}$, and admittance matrices $Y$. The generator-supplied power $P_{\rm m,i}$ as well as the electrical power $P_{{\rm e},i}$ are derived from the AC power flow solution. The generator and load dynamic parameters, $H_{i}$ and $D_{i}$, are estimated using the approach described in [77], with the nominal frequency $f_{\rm R}=60$ Hz. The power flow equations are solved numerically using the MATPOWER toolbox [87]. The initial conditions assume that the power system is in a synchronized steady state, i.e., $\dot{\phi}_{i}(0)=0$, $\forall i$, with $\phi_{i}(0)$ determined by the power flow solution.
    We numerically simulated the system over the time interval $t\in[0,100]$s, introducing a random perturbation $\nu_{i}(t)\sim\mathcal{N}(0,\sigma^{2})$ to the system state $\phi_{i}(0)$ every 2s. This setup mimics the transient response of power grids to random fluctuations, such as changes in load demand, generation, or minor faults. Perturbations are added at regular time intervals (rather than irregularly or continuously) to ensure consistent excitation across time, allowing us to assess the performance of SODAs based on the length of the time-series measurements. Strong and weak perturbations corresponds to $\sigma=0.5$ and $\sigma=0.1$ respectively. We assume that the power grid is equipped with phasor measurement units (PMUs) placed on all $N$ buses, providing us with high-resolution, high-frequency data for phase angle $\phi_{i}(t)$, frequency $\dot{\phi}_{i}(t)$, line voltage $V_{i}(t)$, and electrical power $P_{\rm e,i}(t)$ and $Q_{\rm e,i}(t)$. While this assumption simplifies data availability and ensures state observability [88, 89, 90], it may not be entirely practical for all power grids (particularly those with limited infrastructure or cost constraints, as PMUs are often expensive and not yet deployed universally). However, the increasing adoption of smart-grid technologies and improved data acquisition systems are gradually making this level of measurement more feasible in modern grids [72].
    2 CRN Example Details
    2.1 Model definition
    By the structure presented in Section 4.1, CRN models are defined by a stoichiometry matrix $\mathbf{S}$ and reaction rate vector $\mathbf{v}(\mathbf{x})$. Quasi-steady-state approximations allow a reduced system to be represented in terms of only its slow variables $\mathbf{z}$. For completeness, we give $\mathbf{S_{z}}$ and $\mathbf{v}(\mathbf{g}(\mathbf{z}),\mathbf{z})$ for each CRN where $K_{MN}=\frac{k_{-N}+k_{\text{cat},N}}{k_{N}}$.
    CRN1
    We assume that $AE_{1}$ (and by extension, $E_{1}$ due to mass conservation) is at quasi-steady state, yielding:
    $\mathbf{S}_{\mathbf{z}}$ $=\begin{bmatrix}-1\\ 1\end{bmatrix}$ $\mathbf{v}(\mathbf{g}(\mathbf{z}),\mathbf{z})$ $=\begin{bmatrix}\frac{k_{\text{cat},1}E^{\text{tot}}_{1}[A]}{K_{M1}+[A]}\end{bmatrix}$
    CRN2
    We assume that $BE_{2}$ and $BE_{3}$ (and by extension, $E_{2}$ and $E_{3}$) are at quasi-steady state, yielding:
    $\mathbf{S}_{\mathbf{z}}$ $=\begin{bmatrix}-1&-1\\ 1&0\\ 0&1\end{bmatrix}$ $\mathbf{v}(\mathbf{g}(\mathbf{z}),\mathbf{z})$ $=\begin{bmatrix}\frac{k_{\text{cat},2}E^{\text{tot}}_{2}[B]}{K_{M2}+[B]}\\ \frac{k_{\text{cat},3}E^{\text{tot}}_{3}[B]}{K_{M3}+[B]}\end{bmatrix}$
    CRN3
    We assume that $AE_{1}$, $BE_{2}$, $BE_{3}$, and $DE_{4}$ (and by extension, $E_{1}$, $E_{2}$, $E_{3}$ and $E_{4}$) are at quasi-steady state, yielding:
    $\mathbf{S}_{\mathbf{z}}$ $=\begin{bmatrix}-1&0&0&0\\ 1&-1&-1&0\\ 0&1&0&0\\ 0&0&1&-1\\ 0&0&0&1\\ \end{bmatrix}$ $\mathbf{v}(\mathbf{g}(\mathbf{z}),\mathbf{z})$ $=\begin{bmatrix}\frac{k_{\text{cat},1}E^{\text{tot}}_{1}[A]}{K_{M1}+[A]}\\ \frac{k_{\text{cat},2}E^{\text{tot}}_{2}[B]}{K_{M2,f}+[B]+\left(\frac{K_{M2,f}}{K_{M2,r}}\right)[C]}\\ \frac{k_{\text{-2}}E^{\text{tot}}_{2}[C]}{K_{M2,r}+[C]+\left(\frac{K_{M2,r}}{K_{M2,f}}\right)[B]}\\ \frac{k_{\text{cat},3}E^{\text{tot}}_{3}[B]}{K_{M3}+[B]}\\ \frac{k_{\text{cat},4}E^{\text{tot}}_{4}[D]}{K_{M4}+[D]}\end{bmatrix}$
    Note here that due to reversibility of the reaction catalyzed by $E_{2}$ (green arrow; Figure 3a), $v_{2}$ and $v_{3}$ contain a more complex denominator term than their equivalent reactions nested in CRN2.
    2.2 Non-monotonicity in rediscovery
    Figure 8: Non-monotonic scaling of CRN1 rediscovery with increasing resolution: (a) Recovery pattern for Trial 1 of 10 for 10% noise shows initial monotonicity before resolving at approximately 250 data points per initial condition, (b) Recovery pattern for Trial 1 of 10 for 40% noise; while 40% noise was not included in our scaling study, it is demonstrative of how the recovery pattern can change as noise is increased.
    We were interested in guarantees about data quantity and quality for SODAs success, guiding our choice of definition for SODAs’ data requirement. The choice of this requirement as the resolution above which SODAs always succeeded in recovering true algebraic relationships up to a tested total of 1200 data points per initial condition was also motivated by the fact that recovery appears non-monotonic for a given random seed as the resolution increased up until a certain point. Generally, for low resolutions, recovery was possible, but was subject to the stochasticity introduced by noise. As data increased, recovery became less dependent on a single noise instance. This is what our definition aimed to capture. Examples of non-monotonicity in the discovery of CRN1 are shown for various noise in Figure 8.
    2.3 SVD and Pareto front analysis
    Figure 9: SVD and Pareto front analysis for CRN2 and CRN3: (a) SVD and Pareto front for CRN2 with 1000 data points per initial condition, 0% noise, and $p=2$, (b) SVD and Pareto front for CRN3 with 1000 data points per initial condition, 0% noise, and $p=2$.
    The use of SVD and Pareto front analysis was critical to setting the stopping criterion for library refinement (see Section 3). We include this analysis for CRN1 in the main text in Figure 2. The same analysis is included here for CRN2 and CRN3 in Figure 9. This suggested a stopping criterion of three refinements for CRN2 and five refinements for CRN3.
    2.4 CRN1: Discovered vs. True Model
    Figure 10: Results of CRN1 discovery: Estimation of $\frac{d[A]}{dt}$ for all tested initial conditions with learned coefficient for $[AE_{1}]$ applied to simulated data of 1000 data points per initial condition and $p=2$ (a) (39) is the result of discovery at 0% noise, (b) (40) is the result of discovery at 5% noise, (c) (41) is the result of discovery at 15% noise.
    In testing a single seed at 0%, 5%, and 15% noise, after correct identification of algebraic relationships, we find
    $\frac{d[A]}{dt}=-\frac{d[B]}{dt}=-4.99[AE_{1}],$ (39)
    $\frac{d[A]}{dt}=-\frac{d[B]}{dt}=-5.02[AE_{1}],$ (40)
    $\frac{d[A]}{dt}=-\frac{d[B]}{dt}=-5.07[AE_{1}]$ (41)
    respectively after refitting (Figure 10). This is a result sufficiently close to the true coefficient of $5.00$. We note that when recovery failed, its failures were consistent and a small coefficient was regularly learned in addition to the structure in (39)–(41). We may anticipate that the use of information criteria to determine the optimal sparsity of learned models across a variety of sparsity parameters in LASSO may help distinguish and select among model structures that appear similar.
    3 Pendulum: Discovered vs. True Model
    Figure 11: Prediction based on the discovered model using SODAs: (a) pendulum experiment with negligible damping, (b) heavily damped pendulum animation. In both cases, the highlighted region shows the training data. The simulated solution based on the discovered model matches closely with the entirety of the available data showing good prediction capability.
    We used scipy.optimize.curve_fit function for parameter estimation to determine the correct parameters after SODAs discovered the functional form of the ODEs for the single pendulum. For Case 1, the final discovered solution with optimized parameters is as follows:
    $\ddot{\phi}=-82.9\cos\phi.$ (42)
    For Case 2 with non-zero damping, the final discovered solution with optimized parameters is as follows:
    $\ddot{\phi}=-0.47\dot{\phi}-9.73\cos\phi.$ (43)
    In both cases, simulated solutions based on the discovered equations (42) (Figure 11a) and (43) (Figure 11b) closely aligned with the corresponding training data as well as the remaining processed pixel data. Note that in both cases, the discovered equations (42) and (43) are equivalent to (28) under the translation $\phi=\omega+\frac{3\pi}{2}$, for the cases of zero and non-zero damping, respectively. Note that for Case 1, the parameter $-82.9$ corresponds to $-\frac{g_{\text{modified}}}{l_{\text{modified}}}$, where $g_{\text{modified}}$ represents the modified acceleration due to gravity, and $l_{\text{modified}}$ is the scaled length of the pendulum. Scaling the pixel data involved enforcing $l_{\text{modified}}=1$, which altered the length of the pendulum while preserving the time period. This adjustment resulted in a modified acceleration due to gravity, $g_{\text{modified}}$, in the classical sense. However, for Case 2, since the original pendulum animation depicted a pendulum of length $1$, we were able to obtain a parameter very close to the actual acceleration due to gravity $g$.

### v2

@@ 4.3 Example 3: Non-linear and Chaotic Pendulums
We used a Pareto front of the natural log of the condition number of the library to determine the stopping criterion. In Case 3, we stopped after two iterations as no further improvement in the condition number was observed from the third refinement onward (Figure 5d), implying the existence of two algebraic constraints. We also conducted a comparative study of the data requirement for Cases 2 and 3 as a function of maximum degree of the monomials in the candidate library (Figure 5c). The results indicate that the data requirement, in terms of the number of frames per video segment, stayed within manageable limits. Interestingly, the double pendulum required fewer frames per video segment for successful recovery of the algebraic equations compared to Case 2. This is because the chaotic nature of the pendulum, combined with a 3-minute-long video, provided a good sampling of the state space. We also shuffled the data points along the time axis only for Case 3 to ensure a good representation of the dynamics, given the longer length of the video segment. These factors contributed to the reduced data requirement in terms of the number of frames needed for successful recovery in Case 3. An alternative approach to reduce the number of frames required would be to conduct experiments with more informative initial conditions, which is closely related to experimental design [Seltman [67], Versyck et al. [78]].
@@ 5 Discussion
There are several limitations and challenges worth pursuing in future work. One significant limitation of SODAs is the necessity to observe or measure all the state variables. In many physical systems, measuring all state variables can be challenging due to cost or physical infeasibility. Previous methods such as implicit-SINDy [Mangan et al. [42]] and SINDy-PI [Kaheman et al. [33]] required no measurements of the algebraic states, but also assumed that all states were differential variables. Extending SODAs to discover DAEs even in the presence of unobserved states, potentially through recent strategies [Ribera et al. [60], Bakarji et al. [5], Stepaniants et al. [72]], would enable the identification of algebraic and differential variables and their equations when only a subset of states are measured. However, recovery in this setting will be even more challenging due to the non-convexity of sparse-optimization with hidden states as well as the potential lack of analytical identifiability in nonlinear DAE models [Montanari et al. [51]].
@@
One aspect of DAE discovery we did not yet explore is how SODAs performs when "fast" states are not fully in quasi-steady state. In realistic experimental sampling of systems with fast-slow variables, some transient dynamics of the fast variables may be sampled before they reach quasi-steady state. In this work, we used the reduced ODE formulation (10) of our CRNs to simulate the time series used for equation re-discovery, ensuring that the slow-variables were in quasi-steady state and thus, there was no error in the algebraic relationships. Future work is needed to investigate how SODAs performs as a function of sampling transient vs quasi-steady-state regimes for fast variables. An interesting avenue would be in mimicking the hybrid-SINDY approach [Mangan et al. [40], Li et al. [37]] in which a different model is selected when variables are in quasi-steady state.
Additionally, the discovery of conservation laws for the CRNs could be explored further. Because the total substrate, $M^{tot}$, varied between simulations, identification of the substrate conservation relationship was not possible in our setup, as a single learned $M^{tot}$ cannot fit all simulations. The enzyme conservation relationship could be recovered, as $E_{\text{tot}}$ was held constant between simulations. Inclusion of non-zero initial conditions in $\Theta$ would allow for the discovery of substrate conservation, a close analog to the inclusion of suspected bifurcation parameters in $\Theta$ in the original formulation of SINDy [Brunton et al. [17]] and following work. We chose here to exclude initial conditions from $\Theta$ because when any two (or more) initial conditions are held constant across simulations, they will be in perfect correlation, confounding one another and limiting the interpretability of any resulting algebraic relationships. Additionally, discovery of conservation relationships of this nature up to a constant (e.g., relationships mirroring $\frac{d}{dt}([AE_{1}]+[E_{1}])=0$) was considered, it was not included algorithmically. Inclusion of state derivatives in $\Theta$ would allow for the discovery of relationships of this type, but would require the smoothing of noisy derivatives. We suspect integration with a weak approach [46] may be more appropriate for conservation laws up to a constant.
@@
There is a more general data limitation challenge inherent to sampling dynamical systems: sampling is often determined by the trajectories of the dynamics themselves which induces correlation between library terms. In cases where the available data does not adequately differentiate between lower-order and higher-order terms, strong correlations separate from the true algebraic relationships can occur. In Example 3, we showed how the data requirement increases as the degree of the monomial terms in the library increases to 5. This is attributed to the high correlation between higher-degree monomials in the library. This high correlation between library terms makes it challenging to separately identify their contributions to the dynamics of the system. This issue is not exclusive to SODAs; it is an inherent challenge in all SINDy-type methods that rely on sparse regression with function libraries. This challenge can be overcome by implementing improved data collection strategies or experimental design techniques [Seltman [67], Versyck et al. [78]] that help with distinguishing different candidate library terms. We aim to develop methods to address these challenges in our future work. Finally, this study has focused on DAEs involving only ordinary derivatives, an exciting avenue for future research will be to extend SODAs to DAEs involving partial derivatives.
Data Accessibility: All code and data used for numerical experiments in this paper are available at
https://github.com/mjayadharan/DAE-FINDER_dev/tree/main/Examples.
The DaeFinder package is also available through Python Package Index (PyPI) at
https://pypi.org/project/DaeFinder
@@
Authors’ contributions: M.J. developed the SODAs algorithm, implemented the DaeFinder package, and conducted all experiments (except for Example 1). C.C. contributed to the DaeFinder package and designed and conducted the numerical experiments in Example 1. A.N.M. designed the test cases and assisted in the numerical experiments in Example 2. N.M.M. designed and supervised the research. All authors edited the manuscript and approved the final draft for publication.
Competing Interests: We declare we have no competing interests.
@@
Funding: This material is based on work supported by several funding agencies. N. M. M. and M. J. were supported by the U.S. Department of Energy, Office of Science, Office of Advanced Scientific Computing Research, under Award Number DE-SC0024253 and Army Contracting Command under Award Number W52P1J-21-9-3023. C. C. was supported by the National Science Foundation under Grant No. DGS-2021900. Any opinions, findings, and conclusions or recommendations expressed in this material are those of the author(s) and do not necessarily reflect the views of the Department of Energy, Army Contracting Command, or National Science Foundation.
@@
Acknowledgments: We thank Finn Hagerty for his assistance in running simulations for the CRN example and related discussions.